You spend 40$ on a meal for you and your friends.
Graph the equation 2y+5x=40 where x is the number of sandwiches purchased and y is the number of beverages purchased.
Intercept the intercepts.

Found 3 solutions by Alan3354, josgarithmetic, ikleyn:
Answer by Alan3354(69443)

You spend 40$ on a meal for you and your friends.
Graph the equation 2y+5x=40 where x is the number of sandwiches purchased and y is the number of beverages purchased.
--------------
You graph it. I don't need the practice.
============
Intercept the intercepts.
What?

Answer by josgarithmetic(39617)

---------------------------------------------------------------
You spend 40$ on a meal for you and your friends.
---------------------------------------------------------------

What do you want with that piece of information?


------------------------------------------------------------------
Graph the equation 2y+5x=40 where x is the number of sandwiches purchased and y is the number of beverages purchased.
----------------------------------------------------------------

Expression on the left is cost. Value on the rightside is a specific cost.
-
If x is 0, then 2y=40
y=20
-
If y is 0 , then 5x=40
x=8
-
PLOT THE POINTS (0,20) and (8,0).
Draw the line through the two points.
That will be your graph.

You would probably be most interested in the natural number values on the line.

Answer by ikleyn(52781)

.

As this problem is worded, printed, formulated, posted and presented,

its right place is in the GARBAGE BOX --- not in this forum.